### v1

# MRAG: A Modular Retrieval Framework for Time-Sensitive Question Answering

Abstract
Understanding temporal relations and answering time-sensitive questions is crucial yet a challenging task for question-answering systems powered by large language models (LLMs). Existing approaches either update the parametric knowledge of LLMs with new facts, which is resource-intensive and often impractical, or integrate LLMs with external knowledge retrieval (i.e., retrieval-augmented generation). However, off-the-shelf retrievers often struggle to identify relevant documents that require intensive temporal reasoning. To systematically study time-sensitive question answering, we introduce the TempRAGEval benchmark, which repurposes existing datasets by incorporating temporal perturbations and gold evidence labels. As anticipated, all existing retrieval methods struggle with these temporal reasoning-intensive questions. We further propose Modular Retrieval (MRAG), a trainless framework that includes three modules: (1) Question Processing that decomposes question into a main content and a temporal constraint; (2) Retrieval and Summarization that retrieves evidence and uses LLMs to summarize according to the main content; (3) Semantic-Temporal Hybrid Ranking that scores each evidence summarization based on both semantic and temporal relevance. On TempRAGEval, MRAG significantly outperforms baseline retrievers in retrieval performance, leading to further improvements in final answer accuracy.
affil0affil0affiliationtext: Nanyang Technological University  Alibaba-NTU JRI  NYU Shanghai
Zhejiang University  Center for Data Science, New York University
[…]
MRAG: A Modular Retrieval Framework for
Time-Sensitive Question Answering
1 Introduction
Figure 1: A time-sensitive question example that requires temporal reasoning (as of 6 May 2021 $\rightarrow$ 2019 - 2022) to both retrieve documents and generate answers. State-of-the-art retrieval systems struggle to conduct in-depth reasoning to identify relevant documents. We provide a new diagnostic benchmark TempRAGEval, and propose a new modular framework to tackle this challenge.
Facts are constantly evolving in our ever-changing world. This dynamic nature highlights the need for natural language processing (NLP) systems capable of adapting to new information (Liška et al., 2022; Zhang et al., 2024; Kasai et al., 2024) and providing accurate responses to time-sensitive questions (Chen et al., 2021; Chu et al., 2024). For instance, a common query like “Who is the UK Prime Minister?” sees the answer transition from “Rishi Sunak” to “Keir Starmer” in 2024 (Figure 1).
With developments of large language models (LLMs), existing approaches rely on the parametric knowledge of LLMs to answer time-sensitive questions directly, and constantly update the parametric knowledge on new facts (Rozner et al., 2024; Wu et al., 2024b; Wang et al., 2024). However, updating LLM parameters are often resource-intensive. Moreover, recent studies reveal that even after updates, LLMs often struggle to accurately reflect the latest facts and their associated facts (Cohen et al., 2024). An alternative line of research explores Retrieval-Augmented Generation (RAG), which integrates LLMs with external knowledge (e.g., Wikipedia) through information retrieval (Izacard et al., 2020). While RAG allows for the incorporation of new facts with minimal effort, its performance heavily relies on off-the-shelf retrieval systems, which are often limited to keywords or semantic matching. Time-sensitive questions, however, often require intensive temporal reasoning to identify relevant documents, i.e., reasoning-intensive retrieval (Su et al., 2024a). For example, in Figure 1, retrievers should infer the date “24 July 2019” as relevant to the constraint “as of 6 May 2021”, rather than only match with date like “May 2021”. Despite widespread recognition of the temporal reasoning-intensive retrieval challenge, there remains a lack of measurements and approaches. We fill that lacuna.
We begin by conducting a diagnostic evaluation of existing retrieval approaches for temporal reasoning-intensive retrieval. Following the idea of systematic evaluation with contrast set (Gardner et al., 2020), we repurpose two existing datasets, TimeQA (Chen et al., 2021) and SituatedQA (Zhang and Choi, 2021), to introduce the Temporal QA for RAG Evaluation benchmark (TempRAGEval). We manually augment the test questions with temporal perturbations (e.g., modifying the time period). In addition, we annotate gold evidence on Wikipedia for more accurate retrieval evaluation. As expected, current retrieval methods exhibit limited temporal reasoning capabilities, especially on temporal perturbed questions.
To address time-sensitive questions, we propose a training-free Modular Retrieval framework (MRAG) to enhance temporal reasoning-intensive retrieval. MRAG contains three key modules: (1) Question Processing, which decomposes each question into a main content and a temporal constraint; (2) Retrieval and Summarization, which adopts off-the-shelf retrievers to find evidence passages based on the main content, then instructs LLM to condense or split relevant passages into query-related sentences; (3) Semantic-Temporal Hybrid Ranking, which ranks each evidence sentence using a combination of a semantic score measuring semantic similarity, and a temporal score, a novel symbolic component that assesses temporal relevance to the query’s temporal constraint.
Compared to the classic retrieve-then-rerank systems, MRAG offers several advancements: (1) It disentangles relevance-based retrieval from temporal reasoning through question preprocessing, avoiding distractions of off-the-shelf retrievers; (2) It introduces a temporal scoring mechanism that evaluates whether the temporal information in the evidence (e.g.,2019 – 2022”) satisfies the query’s temporal relations (e.g.,before 2023”); (3) The modular framework enhances transparency, enabling users to easily identify which component leads to an incorrect answer.
On TempRAGEval, our proposed MRAG framework achieves substantial improvements in performance, with 9.3% top-1 answer recall and 11% top-1 evidence recall. We also incorporate state-of-the-art (SOTA) answer generators (Asai et al., 2024; Yan et al., 2024; Wei et al., 2022), and demonstrate that the improvements in retrieval from MRAG propagate to enhanced final QA accuracy, with 4.5% for both exact match and F1. Detailed case studies further confirms MRAG’s robustness to temporal perturbations qualitatively.
Our contributions can be summarized as follows.
  • We introduce TempRAGEval, a time-sensitive QA benchmark dataset to diagnostically evaluate each component of existing retrieval-augmented generation systems.
  • We propose MRAG, a new modular framework to identify evidence through separating relevance-based retrieval and temporal reasoning.
  • On TempRAGEval, MRAG significantly outperforms all baseline retrieval systems, and the improvements further lead to better answer predictions.
2 Background
In this section, we first define the time-sensitive question answering task (§2.1), and then introduce the baseline retrieval-augmented generation QA based systems (§2.2).
2.1 Time-sensitive Question Answering
Our task is to predict an answer $y$, to a time-sensitive question $q$, in which the input question $q$ also includes temporal relations $t_{1},...,t_{n}$. [^1] To simplify our setting, we make sure each question has a valid answer.
With advancements in LLMs, a straightforward approach is to directly ask these questions to an off-the-shelf LLM (e.g., LLama 3.1), hoping the relevant facts are captured in its parametric knowledge. However, this is often not the case for time-sensitive questions. A significant portion of such questions are recent events (e.g., “Who is the UK Prime Minister in November 2024?”), which typically fall beyond the knowledge cutoff for training LLMs (e.g., December 2023 for LLama 3.1).
One potential solution is to constantly update the parameters of LLMs with new facts, i.e., knowledge editing (Wu et al., 2024b; Wang et al., 2024). However, this approach is highly resource-intensive and often impractical for very large LLMs (e.g., GPT-4o). Moreover, knowledge editing is not always effective. Recent studies reveal that even after parameter updates, LLMs often fail to accurately answer questions about new facts (Liška et al., 2022; Rozner et al., 2024; Wang et al., 2024), let alone related or associated facts (Zhong et al., 2023; Wu et al., 2024b, c).
Therefore, our study adopts an alternative approach that leverages external text collections as a knowledge source, integrating relevant information into existing LLMs to answer questions (detailed in §2.2). While we use Wikipedia222we use December 2021 Wikipedia dump, with 33.1M text chunks as the text collection $D$, our approach is broadly applicable to any corpus, e.g., ClueWeb (Overwijk et al., 2022).
2.2 Retrieval-Augmented Generation
The goal of RAG is to address the limitations in the parametric knowledge of LLMs (Izacard et al., 2020). RAG involves a retrieval system to find relevant evidence from a large corpus and a LLM to generate a response based on the evidence (Wu et al., 2024a).
Passage retrieval and reranking.
Retrieval methods are typically categorized into sparse retrieval and dense retrieval. Sparse retrieval methods, like TF-IDF (Das and Chakraborty, 2018) and BM25 (MacAvaney et al., 2020), rely on lexical matching. In contrast, dense retrieval models (Karpukhin et al., 2020; Zhao et al., 2021; Thakur et al., 2021; Izacard et al., 2021) use LLMs (e.g., BERT) encode the question $q$ and the passage $p$ separately using two independent encoders (i.e., bi-encoders). These models learn a scoring function (e.g., dot product) between question and passage vectors to capture semantic similarities:
$f(q,p)=\text{sim}(\text{Enc}_{Q}(q),\text{Enc}_{P}(p)),p\in\mathcal{D}.$ (1)
Dense retrieval models are highly scalable, since passages can be encoded offline, and are efficiently retrieved over maximum inner product search (MIPS) with the question (Shrivastava and Li, 2014).
However, bi-encoder models may lack the ability to capture fine-grained interactions between the query and passage. A common optional [^3] Note that reranking is not always adopted, as it adds additional computational cost. approach is to have another cross-encoder model to rerank top retrieved passages. Cross-encoder models (Khattab and Zaharia, 2020; Wang et al., 2020; Gemma et al., 2024) jointly encode the query $q$ and the passage $p$ together by concatenating them as input into a single model as:
$f(q,p)=\text{sim}(\text{Enc}([q;p])),p\in\mathcal{D}.$ (2)
Answer generation.
The answer generation component takes the question and support passages as input to generate the answer. A widely used approach is Fusion-in-Decoder (Izacard et al., 2020), which fine-tunes an encoder-decoder LLM (e.g., T5 (Raffel et al., 2020)). FiD independently encodes each passage and question, then concatenates their representations into the decoder, as formulated below:
$y=\text{Dec}([\text{Enc}([q;p_{1}]);\dots;\text{Enc}([q;p_{k}])]),p_{k}\in\mathcal{D}.$ (3)
With decoder-only LLMs (OpenAI et al., 2024; Dubey et al., 2024), the support passages and the question are concatenated into a single sequence.
To unlock the reasoning capabilities of LLMs, Chain-of-Thought (CoT) prompting (Wei et al., 2022) introduces intermediate reasoning steps within prompt examples, significantly improving performance on various reasoning tasks. Self-RAG (Asai et al., 2024) predicts answers for each passage independently, and then adopts self-reflect mechanisms (Yao et al., 2022) into prompts to select the best answer. Self-RAG achieve SOTA results on multiple knowledge-intensive QA datasets.
3 TempRAGEval Benchmark
In this section, we first present existing time-sensitive QA datasets (§3.1), then introduce our diagnostic benchmark dataset TempRAGEval3.2), finally we evaluate existing retrieval approaches on TempRAGEval3.3).
3.1 Existing Time-Sensitive QA Datasets
Existing time-sensitive QA datasets often feature questions where timestamps are closely tied to their answers, as illustrated in Figure 1. Consequently, retrievers relying on surface-form date matching tend to perform well on these test sets, limiting the evaluation of true temporal reasoning capabilities. To systematically evaluate time-sensitive question answering, we establish the following criteria: (1) questions with varying difficulty levels, including both common and rare entities; (2) manual evidence annotation to enable retrieval evaluation; and (3) temporal perturbations as contrast sets (Gardner et al., 2020). Since no existing datasets provide both temporal perturbations and evidence annotations, we primarily adapt from the following two datasets:
  • SituatedQA (Zhang and Choi, 2021) is a open-domain QA dataset where the correct answer to a question is based on a temporal context, specifically of the type “as of”. SituatedQA sources questions from exiting QA datasets such as NQ-Open (Kwiatkowski et al., 2019) and WebQuestions (Berant et al., 2013), which primarily focus on popular entities.
  • TimeQA (Chen et al., 2021) is a time-sensitive QA dataset in which question-answer pairs are synthetically generated from time-evolving WikiData facts using templates. The dataset includes a wide range of reasoning types and difficulty levels, with a notable emphasis on rare entities.
Other time-sensitive QA datasets, StreamingQA Liška et al. (2022), TempLAMA (Dhingra et al., 2022), and concurrent dataset ComplexTQA (Gruber et al., 2024), can also serve as alternatives. Temporal reasoning benchmarks such as Wei et al. (2023); Chu et al. (2024); Wang and Zhao (2024); Tan et al. (2023) fall outside the scope of our study.
3.2 TempRAGEval Construction
Dataset # Eval. Evid. Natu. Pert.
ComplexTQA 10M
StreamingQA 40K $\checkmark$ $\checkmark$
TempLAMA 35K
SituatedQA 2K $\checkmark$
TimeQA 3K
MenatQA 2K
TempRAGEval 1K $\checkmark$ $\checkmark$ $\checkmark$
Table 1: Comparison of temporal QA datasets. TempRAGEval is featured by manual evidence annotations, human-written question (i.e., Naturalness), and perturbations.
We create TempRAGEval, a time-sensitive QA benchmark for rigorously evaluating temporal reasoning in both retrieval and answer generation.
Perturbed question-answer pair generation.
We source questions from TimeQA (Chen et al., 2021) and SituatedQA (Zhang and Choi, 2021) datasets. For each question, we introduce perturbations to the original temporal constraints, e.g.,as of 2019”, with complex variations such as e.g.,between 2020 and 2021”, “the last before 2020”, or “as of 6 May 2021”. We guarantee the answer (e.g.,Boris Johnson”) remains consistent across the perturbed questions.
Evidence annotation.
To better evaluate the performance of retrieval systems, we supplement question-answer pairs with up to two annotated gold evidence passages. Specifically, for each question, we manually review top-100 passages retrieved by the leading retriever, Contriever (Izacard et al., 2021), and the best reranker, Gemma (Li et al., 2023). If these retrieval methods fail to identify relevant passages within top-100 passages, we manually search Wikipedia pages related to the query entities to locate the gold evidence (around 12.7% of original questions). In total, we have annotated 1,000 test examples with both original and perturbed questions, along with human-annotated evidence. TempRAGEval-TimeQA (500 examples) focus on popular entities, and TempRAGEval-TimeQA (500 examples) on rare entities.
3.3 Preliminary Evaluation on TempRAGEval
In TempRAGEval, we first evaluate the performance on SOTA retrieval systems as a sanity check.
Experimental Setup.
We follow the popular retrieve-then-rerank pipeline, using the dense retriever Contriever (Izacard et al., 2021) and the LLM-based reranker Gemma (Gemma et al., 2024). The retriever finds top 1,000 passages, and among them, the reranker reorders the top 100 passages. We use two evaluation metrics: Answer Recall (AR@k) that measures the proportion of samples where at least one answer appears within the top-k retrieved passages, and Gold Evidence Recall (ER@k) that assesses the percentage of samples where at least one gold evidence document is included in the top-k passages.
Performance degradation on perturbed questions.
As shown in Figure 2, we observe a significant degradation in retrieval performance caused by temporal perturbations. For instance, for the Gemma baseline, the top-1 answer recall and evidence recall drop from 85.8% to 54.7% and from 45.0% to 20.3% on TempRAGEval-SituatedQA. This is because the perturbed temporal constraints avoid matching between timestamps in the questions and the passages. Consequently, retrievers must conduct in-depth temporal reasoning to identify the relevant passages.
We further conduct a controlled experiment to reveal the temporal reasoning capabilities of existing retrieval methods. Specifically, we compute the similarity scores for query-evidence pairs by varying the temporal relation in the query, e.g.,before”, “after”, and “as of”, and the timestamp in the evidence, e.g., from “1958” to “1965”. Experiments confirm that all methods prioritize matching exact dates indicating a shortcut for temporal reasoning in retrieval. We present full results in Figure 5 in Appendix.
Figure 2: The retrieval performance degradation of the Gemma baseline on TempRAGEval-SituatedQA, comparing original and perturbed questions (see TempRAGEval-TimeQA in Appendix C).
4 MRAG: Modular Retrieval
Figure 3: An overview of the MRAG framework, consisting of three key modules: question processing, retrieval and summarization, and semantic-temporal hybrid ranking. The question processing module separates each query into the main content (i.e., MC) and the temporal constraint (i.e., TC). The retrieval and summarization module finds the most relevant evidence based on the main content and summarizes or splits these evidence into fine-grained sentences. The hybrid ranking module combines symbolic temporal scoring and dense embedding-based semantic scoring at a fine-grained level to determine the final evidence ranking.
Motivated by the performance degradation of the existing retrieval methods in the preliminary evaluation, we propose a Modular Retrieval (MRAG) framework (as shown in Figure 3) to enhance temporal reasoning-intensive retrieval. At a high level, MRAG disentangles relevance-based retrieval from temporal reasoning, leveraging a dense embedding model for semantic scoring and a set of symbolic heuristics for temporal scoring. Specifically, MRAG has three key modules: question processing, retrieval and summarization, and semantic-temporal hybrid ranking.
Question processing.
We prompt LLMs to decompose each time-sensitive question into a main content (MC) and a temporal constraint (TC). This approach disentangles temporal relevance from semantic relevance: MC measures the semantic relevance of the evidence, while TC determines its temporal relevance.
Retrieval and summarization.
We apply off-the-shelf retrievers (e.g., Contriever) to find relevant passages to MC in Wikipedia. Then we adopt a cross-encoder model to rerank the passages with high semantic similarity to MC.
It is common for a passage to contain multiple pieces of temporal information, most of which are unrelated to the question and can distract the temporal scoring component introduced next. For example, the relevant passage in Figure 3 includes the sentence, “it was filmed in 2017”, which satisfies the TC but is irrelevant. Therefore, we initially considered splitting passages into individual sentences to eliminate temporal distractors. However, as shown in Figure 3, critical information from the most relevant passages—such as “America’s Next Top Model”, “The winner of the competition”, and ”January 9, 2018”—is often scattered across different sentences. Relying solely on sentence splitting would miss key details. To overcome this challenge, we instead adopt LLMs to summarize the top-k passages into a single sentence to condense relevant phrases and temporal information. [^4] While using LLMs to summarize passages can introduce hallucinations, we mitigate this by summarizing only the top-k passages and simply splitting the remaining passages into sentences. An analysis of this approach is provided in §6.1.
Semantic-temporal hybrid ranking.
We rerank each sentence (summarized from a LLM or segmented from the original passage) with two distinct scores: a semantic score and a temporal score. The semantic score is calculated from the similarity between the evidence sentence and MC. For temporal score, we first extract temporal information in the evidence sentence, then we apply symbolic scoring functions [^5] We include details of pre-defined temporal scoring functions in Appendix D. to verify whether the temporal information in the evidence satisfies TC. For example, the evidence summary timestamp “2018” satisfies TC “as of 2021” in Figure 3. The final score for each sentence is the product of the semantic score and the temporal score. We select the passages that contains the highest-scoring sentences.
5 Experiments
In this section, we evaluate MRAG and baseline systems on TempRAGEval.
5.1 Experimental Setup
Baselines.
For retrieval, we include BM25, Contriever, and a hybrid method (Jedidi et al., 2024). Reranking methods include ELECTRA (Clark et al., 2020), MiniLM (Wang et al., 2020), Jina(Jina, 2024), BGE (Xiao et al., 2023), NV-Embed (Lee et al., 2024), and Gemma (Gemma et al., 2024). We follow state-of-the-art answer generation approaches based on prompting LLMs (§2.2). We evaluate four approaches, Direct Prompt that adopts question-answer pairs as few-shot examples; Direct CoT that adds rationals into prompts (Wei et al., 2022); RAG-Concat, where passages are concatenated into a LLM; and Self-RAG, which processes each passage independently and selects the best answer (Asai et al., 2024).
Metrics.
We use the same setup in §3.3 for retrieval evaluation. For answer evaluation, we use Exact Match (EM) that measures the exact match to the gold answer, and F1 score (F1) that measures the word overlap to the gold answer.
Implementation details.
Due to limited budget, we evaluate GPT-4o mini (OpenAI et al., 2024) with direct prompting, and three open-source LLMs: TIMO, a LLaMA2-13B model fine-tuned for temporal reasoning (Su et al., 2024b), and two general-purpose models, Llama3.1-8B-Instruct and Llama3.1-70B-Instruct Dubey et al. (2024). We use 10 examples in prompts.
Method TempRAGEval-TimeQA TempRAGEval-SituatedQA
AR @ ER @ AR @ ER @
1st 2nd # QFS 1 5 1 5 1 5 1 5
BM25 - - 17.5 39.0 4.2 14.1 27.6 58.2 6.8 18.4
Cont. - - 18.8 49.9 9.6 28.7 22.6 51.1 6.8 17.1
Hybrid - - 18.8 51.2 9.6 28.1 22.6 55.8 6.8 19.7
Cont. ELECTRA - 40.1 76.9 21.8 58.6 35.5 71.3 15.3 37.1
Cont. MiniLM - 34.0 76.1 16.2 57.3 36.8 73.4 20.0 40.3
Cont. Jina - 42.4 77.2 23.6 58.6 47.9 78.4 19.5 41.1
Cont. BGE - 40.3 80.9 23.3 61.3 36.3 74.2 14.5 35.0
Cont. NV-Embed - 49.9 81.2 33.4 62.9 47.4 81.3 23.4 46.1
Cont. Gemma - 46.7 82.5 26.0 66.6 54.7 82.6 20.3 45.3
Cont. MRAG - 57.6 89.4 32.4 73.5 61.1 88.2 27.4 56.3
Cont. MRAG 5 58.6 90.0 37.1 74.3 61.3 89.0 31.1 59.2
Cont. MRAG 10 56.0 88.1 35.5 73.2 62.1 87.9 30.8 57.9
Table 2: The answer recall (AR@k) and gold evidence recall (ER@k) of each retrieval method on perturbed temporal queries in TimeQA and SituatedQA subsets of TempRAGEval. 1st means the first-stage retrieving method; 2nd means the second-stage reranking method; # QFS means the number of top passages to be summarized. Bold numbers indicate the best performance. We include complete results in Appendix E.
5.2 Main Results
Method TempRAGEval-TimeQA TempRAGEval-SituatedQA
# Docs EM F1 # Docs EM F1
GPT4o-mini
Direct Prompt - 19.6 30.6 - 54.2 58.6
TIMO
Direct Prompt - 16.2 24.8 - 50.6 53.1
Direct CoT - 15.8 28.2 - 49.4 53.9
RAG-Concat 3 43.4 55.2 3 55.8 58.1
MRAG-Concat 3 48.2 57.2 3 61.4 63.6
Self-MRAG 3 44.6 54.9 3 62.4 64.5
Llama3.1-8B-Instruct
Direct Prompt - 16.0 23.9 - 42.8 45.0
Direct CoT - 16.8 27.8 - 49.6 54.5
RAG-Concat 5 44.0 52.8 5 60.0 62.7
MRAG-Concat 5 49.2 59.2 5 65.8 68.0
Self-MRAG 5 54.2 65.6 5 66.4 68.2
Llama3.1-70B-Instruct
Direct Prompt - 31.0 42.3 - 59.0 62.1
Direct CoT - 33.2 45.8 - 69.0 72.6
RAG-Concat 5 54.4 63.2 20 67.0 69.8
MRAG-Concat 5 58.0 68.4 20 69.2 72.5
Self-MRAG 5 61.2 75.3 20 72.2 76.0
Table 3: End-to-end QA comparison for various generation strategies and LLMs on TempRAGEval. Bold numbers indicate the best performance the each backbone LLM. The second best is underlined.
MRAG enhances retrieval performance for time-sensitive questions.
According to Table 2, MRAG significantly outperform all retrieve then rerank baselines, which highlight the superior temporal reasoning capabilities. For example, MRAG improves the best baselines Contriver + Gemma significantly, with 7.7% top-5 evidence recall in TempRAGEval-TimeQA and 13.9% top-5 evidence recall in TempRAGEval-SituatedQA.
Retrieval augmentation improves time-sensitive QA performance.
According to Table 3, we observe that LLMs relying solely on their parametric knowledge struggle to accurately answer time-sensitive questions, with limited QA accuracy. Incorporating retrieval-augmented generation significantly improves QA accuracy. Notably, we observe larger improvements on TempRAGEval-TimeQA, which primarily focuses on less frequent entities that pose greater challenges to the parametric knowledge of LLMs (Kandpal et al., 2023).
Enhanced retrieval contributes to improved time-sensitive QA performance.
As shown in Table 3, MRAG outperforms baseline RAG approaches in QA accuracy, for instance 49.2% EM (MRAG) over 44.0% EM (RAG) in TempRAGEval-TimeQA for Llama3.1-8B. This highlights the role of enhanced retrieval in improving accuracy of the time-sensitive QA task. Incorporating a self-reflection strategy improves performance for Llama3.1 models but not for TIMO. We observe that TIMO makes more errors in self-reflection, likely due to the limited reasoning capacity of its backbone model, Llama 2.
6 Analysis
In this part, we provide detailed analysis on how each component of MRAG contributes to the overall improvement.
6.1 Ablation Study
The impact of the numbers of passages for summarization.
Our LLM based summarization removes irrelevant temporal information but may also introduce hallucinations. We conduct a parametric study to evaluate the impact of the number of passages used for summarization. As shown at the bottom of Table 2, summarizing top-five passages achieves the best balance.
The impact of the number of passages used for answer generation in MRAG.
Our experiments show that the optimal number of passages depends on the size of the LLM. TIMO can handle a maximum of 3 passages, while the optimal number for LLaMA 3.1-8B is five, and for LLaMA 3.1-70B, it is twenty. Larger LLMs, with more parameters, process longer contexts better. Full results are presented in LABEL:sec:_concat.
6.2 Human Evaluation
One limitation of the retrieval metrics is that AR@k overestimates performance, as a passage might incidentally contain an answer without directly supporting it. Conversely, ER@k acts as a conservative lower bound, potentially overlooking other relevant but unannotated passages. To address this, we conduct a human evaluation on a random subset of 200 examples to assess actual retrieval performance. We manually review the top 20 passages retrieved by MRAG and Gemma, annotating the first relevant passage. The results confirm that MRAG consistently outperforms Gemma in retrieval accuracy. We present full results in LABEL:sec:_human.
6.3 Case Study
Figure 4: A case study for top-1 passages retrieved by Gemma and MRAG for a question from TempRAGEval-SituatedQA.
We conduct case studies to qualitatively assess the advantages of MRAG over the Gemma baseline. According to Figure 4, the top-1 passage by Gemma matches the query date “1988”, but discusses a father-son record set in 2007. In contrast, MRAG result is about a teammate combination record in the same season, despite the date “1961” differing from the query date “1988”. We include more cases in LABEL:sec:_cases.
7 Other Related Works
LLM embeddings.
Recent research has explored LLM embeddings for retrieval. Some studies focus on distilling or fine-tuning LLM embeddings for reranking tasks, such as Gemma (Gemma et al., 2024) and MiniCPM (Hu et al., 2024). Others aim to develop generalist embedding models capable of performing a wide range of tasks including retrieval and reranking, e.g., gte-Qwen (Yang et al., 2024) and NV-Embed (Lee et al., 2024). These LLM-based methods have demonstrated unprecedented performance in benchmarks such as MTEB (Muennighoff et al., 2023) over Contriver-based dense retrievers, and also on our TempRAGEval benchmark.
Reasoning intensive retrieval.
Existing retrieval benchmarks primarily target keyword-based or semantic-based retrieval. Su et al. (2024a) introduces BRIGHT, a new retrieval benchmark emphasizing intensive reasoning. The benchmark challenges retrievers to perform reasoning across diverse domains, such as analogies and algorithms, to identify relevant documents. We focus on temporal reasoning-intensive retrieval, sharing a similar vision for future advanced retrieval systems.
8 Conclusion
This study focuses on time-sensitive QA, a task that challenges LLM based QA systems. We first present TempRAGEval, a diagnostic benchmark featuring natural questions, evidence annotations, and temporal perturbations. We further propose a training-free MRAG framework, which disentangles relevance-based retrieval from temporal reasoning and introduces a symbolic temporal scoring mechanism. While existing systems struggle on TempRAGEval due to limited temporal reasoning capacities in retrieval, MRAG shows significant improvements. We hope this work advances future research on reasoning-intensive retrieval.
Limitations
There are still some limitations in our work: (1) Our proposed benchmark is designed to evaluate time-sensitive questions with explicit temporal constraints. However, addressing questions with implicit temporal constraints presents a more complex challenge for retrieval systems. Our symbolic approach is not directly applicable to such questions and requires more advanced reasoning mechanisms to bridge this gap. (2) Our benchmark primarily targets time-sensitive questions with a single answer. Future work can extend to predict a list of answers with different temporal constraints. (3) Despite our approach improves temporal reasoning in retrieval, there is still head room. Fine-tuning dense retrieval models on temporal reasoning-intensive retrieval datasets may offer a promising solution for future work.
Appendix A Controlled Experiments
We conduct controlled experiments to investigate the behaviors of retrieval methods on temporally constrained queries, including the bi-encoder retriever Contriever (Izacard et al., 2021), the cross-encoder reranker MiniLM (Wang et al., 2020), and the LLM embedding-based reranker Gemma (Li et al., 2023). As shown in Figure 5, all methods prioritize date-matching, having the highest scores when the query and document share the same year. Besides, the score is unusually high when the query and document share the same month and day but differ in year, e.g., orange triangles in the diagrams.
The Contriever retriever is less sensitive to document dates than the MiniLM reranker. Both methods exhibit similar trends across varying temporal relations, indicating their inability to differentiate effectively between different relations, e.g., “before” and “after”. Notably, documents without specific dates receive unusually low scores, even lower than those with irrelevant dates, e.g., orange dash lines in the diagrams.
The LLM embedding based reranker Gemma exhibits stronger temporal reasoning capabilities. For the “after” relation, documents with dates later than the query date are assigned relatively high and consistent scores. So all temporally relevant documents will be retained. However, for “before” and “as of”, despite their temporal relevance, documents with earlier dates fail to achieve sufficiently high similarity scores, potentially leading to their exclusion from the retrieval process.
In summary, existing retrieval methods demonstrate limited temporal reasoning capabilities. The LLM embedding-based method shows better performance than others.
Figure 5: Similarity scores of query-document pairs by varying the temporal relation in the query and the date in the document.
Appendix B Implementation Details
We conduct empirical evaluations for MRAG and SOTA retrieving-and-reranking systems on TempRAGEval. MRAG consists of functional modules, which can be based on algorithms, models, or prompting methods. In implementation, algorithm based modules include question normalization, keyword ranking, time extraction, and semantic-temporal hybrid ranking. Model based modules are retrieving, semantic ranking, sentence tokenization. To ensure fair comparison, we use the same retriever model, i.e., Contriever (Izacard et al., 2021), as the first stage method for MRAG and two-stage systems. We use Gemma embeddings (Li et al., 2023) as the main tool to measure semantic similarity for passages and sentences in MRAG. NLTK package is used for sentence tokenization (Bird and Loper, 2004). LLM prompting based modules are keyword extraction, query-focused summarization. As shown in Appendix E, we have tested Llama3.1-8B-Instruct and Llama3.1-70B-Instruct models (Dubey et al., 2024) for LLM prompting based modules. Detailed prompts are listed in LABEL:sec:_prompts. The evaluation metrics are computed based on retrieved passages not sentences.
Appendix C Retrieval Performance Degradation from Perturbations
We compare the retrieval performance between original queries and perturbed queries using the same baseline retrieval system (i.e., Contriever retriever and Gemma reranker). For both the TimeQA and SituatedQA subsets, the perturbed questions significantly increase the difficulty of retrieving relevant documents, particularly when evaluating the top-1 and top-5 ranked documents. This suggests that the introduction of perturbations introduces greater complexity. The existing retrieval method has limited temporal reasoning capabilities and is not robust to such variations.
Figure 6: Retrieval performance difference between original queries and perturbed queries in TempRAGEval subsets for the baseline Gemma retrieval.
Appendix D Spline Functions for Temporal Scoring
In MRAG, the temporal relevance is determined symbolically using temporal constraints in query, extracted dates from sentences, and pre-defined scoring functions. The temporal constraints are categorized by implicit conditions, i.e., “first” and “last”, and temporal relations, i.e., “before”, “after”, and “between”. For example, the question “Who won the latest game as of 1981?” will be classified as the “last” type and the “before” type. We pre-define spline function templates for each type of temporal constraints, as shown in Figure 7. Each spline function computes a temporal score for a given input date extracted from the sentence. For sentences with multiple dates, the highest temporal score is selected.
Figure 7: Pre-defined spline functions for temporal relevance scoring. The title of each subplot represents the type of query temporal constraint. The horizontal coordinate of each subplot is the date in the document sentence.
Appendix E Complete Retrieval Evaluation Results

### v2

@@
 Abstract
-Understanding temporal relations and answering time-sensitive questions is crucial yet a challenging task for question-answering systems powered by large language models (LLMs).
+Understanding temporal concepts and answering time-sensitive questions is crucial yet a challenging task for question-answering systems powered by large language models (LLMs).
 Existing approaches either update the parametric knowledge of LLMs with new facts, which is resource-intensive and often impractical, or integrate LLMs with external knowledge retrieval (i.e., retrieval-augmented generation).
 However, off-the-shelf retrievers often struggle to identify relevant documents that require intensive temporal reasoning.
-To systematically study time-sensitive question answering, we introduce the TempRAGEval benchmark, which repurposes existing datasets by incorporating temporal perturbations and gold evidence labels.
+To systematically study time-sensitive question answering, we introduce the TempRAGEval benchmark, which repurposes existing datasets by incorporating complex temporal perturbations and gold evidence labels.
 As anticipated, all existing retrieval methods struggle with these temporal reasoning-intensive questions.
-We further propose Modular Retrieval (MRAG), a trainless framework that includes three modules: (1) Question Processing that decomposes question into a main content and a temporal constraint; (2) Retrieval and Summarization that retrieves evidence and uses LLMs to summarize according to the main content; (3) Semantic-Temporal Hybrid Ranking that scores each evidence summarization based on both semantic and temporal relevance.
+We further propose Modular Retrieval (MRAG), a trainless framework that includes three modules: (1) Question Processing that decomposes question into a main content and a temporal constraint; (2) Retrieval and Summarization that retrieves, splits, and summarize evidence passages based on the main content; (3) Semantic-Temporal Hybrid Ranking that scores semantic and temporal relevance separately for each fine-grained evidence.
 On TempRAGEval, MRAG significantly outperforms baseline retrievers in retrieval performance, leading to further improvements in final answer accuracy.
+[^1] Our code and data are available at https://github/anonymous.
 affil0affil0affiliationtext: Nanyang Technological University  Alibaba-NTU JRI  NYU Shanghai
@@
 Facts are constantly evolving in our ever-changing world.
-This dynamic nature highlights the need for natural language processing (NLP) systems capable of adapting to new information (Liška et al., 2022; Zhang et al., 2024; Kasai et al., 2024) and providing accurate responses to time-sensitive questions (Chen et al., 2021; Chu et al., 2024).
-For instance, a common query like “Who is the UK Prime Minister?” sees the answer transition from “Rishi Sunak” to “Keir Starmer” in 2024 (Figure 1).
-With developments of large language models (LLMs), existing approaches rely on the parametric knowledge of LLMs to answer time-sensitive questions directly, and constantly update the parametric knowledge on new facts (Rozner et al., 2024; Wu et al., 2024b; Wang et al., 2024).
+This dynamic nature highlights the need for natural language processing (NLP) systems capable of updating information (Liška et al., 2022; Zhang et al., 2024; Kasai et al., 2024) and providing accurate responses to time-sensitive questions (Chen et al., 2021; Chu et al., 2024).
+For instance, a common query like “Who is the UK Prime Minister?” sees the answer transition from “Boris Johnson” to “Rishi Sunak” in 2022 (Figure 1).
+With developments of large language models (LLMs), existing approaches rely on the parametric knowledge of LLMs to answer time-sensitive questions directly, and constantly update the parametric knowledge on new facts (Rozner et al., 2024; Wu et al., 2024a; Wang et al., 2024).
 However, updating LLM parameters are often resource-intensive.
-Moreover, recent studies reveal that even after updates, LLMs often struggle to accurately reflect the latest facts and their associated facts (Cohen et al., 2024).
 An alternative line of research explores Retrieval-Augmented Generation (RAG), which integrates LLMs with external knowledge (e.g., Wikipedia) through information retrieval (Izacard et al., 2020).
@@
 For example, in Figure 1, retrievers should infer the date “24 July 2019” as relevant to the constraint “as of 6 May 2021”, rather than only match with date like “May 2021”.
-Despite widespread recognition of the temporal reasoning-intensive retrieval challenge, there remains a lack of measurements and approaches.
+Although temporal QA is a well-established domain, there remains a lack of research for temporal reasoning-intensive retrieval systems.
 We fill that lacuna.
@@
 Following the idea of systematic evaluation with contrast set (Gardner et al., 2020), we repurpose two existing datasets, TimeQA (Chen et al., 2021) and SituatedQA (Zhang and Choi, 2021), to introduce the Temporal QA for RAG Evaluation benchmark (TempRAGEval).
-We manually augment the test questions with temporal perturbations (e.g., modifying the time period).
+We manually augment the test questions with complex temporal perturbations (e.g., modifying the time period to avoid textual overlap).
 In addition, we annotate gold evidence on Wikipedia for more accurate retrieval evaluation.
-As expected, current retrieval methods exhibit limited temporal reasoning capabilities, especially on temporal perturbed questions.
-To address time-sensitive questions, we propose a training-free Modular Retrieval framework (MRAG) to enhance temporal reasoning-intensive retrieval.
-MRAG contains three key modules: (1) Question Processing, which decomposes each question into a main content and a temporal constraint; (2) Retrieval and Summarization, which adopts off-the-shelf retrievers to find evidence passages based on the main content, then instructs LLM to condense or split relevant passages into query-related sentences; (3) Semantic-Temporal Hybrid Ranking, which ranks each evidence sentence using a combination of a semantic score measuring semantic similarity, and a temporal score, a novel symbolic component that assesses temporal relevance to the query’s temporal constraint.
-Compared to the classic retrieve-then-rerank systems, MRAG offers several advancements: (1) It disentangles relevance-based retrieval from temporal reasoning through question preprocessing, avoiding distractions of off-the-shelf retrievers; (2) It introduces a temporal scoring mechanism that evaluates whether the temporal information in the evidence (e.g.,2019 – 2022”) satisfies the query’s temporal relations (e.g.,before 2023”); (3) The modular framework enhances transparency, enabling users to easily identify which component leads to an incorrect answer.
+We observe remarkable degradation of retrieval performance, which points out a serious robustness issue for existing retrievers.
+To address this issue, we propose a training-free Modular Retrieval framework (MRAG) to enhance temporal reasoning-intensive retrieval.
+Complementary to recent agentic RAG methods such as R1-Searcher and Search-o1 (Song et al., 2025; Li et al., 2025), MRAG provides an improved retrieval component and reduces the number of invocations for retrieval when handling complex temporal questions.
+Specifically, MRAG contains three key modules: (1) Question Processing, which decomposes each question into a main content and a temporal constraint; (2) Retrieval and Summarization, which utilizes off-the-shelf retrievers to find evidence passages based on the main content, segments them into independent sentences, and guides LLMs to condense the most relevant passages into query-specific sentences.(3) Semantic-Temporal Hybrid Ranking, which ranks each evidence sentence using a combination of a semantic score measuring semantic similarity, and a temporal score, a novel symbolic component that assesses temporal relevance to the query’s temporal constraint.
 On TempRAGEval, our proposed MRAG framework achieves substantial improvements in performance, with 9.3% top-1 answer recall and 11% top-1 evidence recall.
@@
 Our contributions can be summarized as follows.
-  • We introduce TempRAGEval, a time-sensitive QA benchmark dataset to diagnostically evaluate each component of existing retrieval-augmented generation systems.
-  • We propose MRAG, a new modular framework to identify evidence through separating relevance-based retrieval and temporal reasoning.
-  • On TempRAGEval, MRAG significantly outperforms all baseline retrieval systems, and the improvements further lead to better answer predictions.
+  • We introduce TempRAGEval, a time-sensitive RAG benchmark to diagnostically evaluate each component of existing retrieval-augmented generation systems.
+  • We propose MRAG, a modular retrieval framework to separately determine semantic and temporal relevance.
+  • On TempRAGEval, MRAG significantly outperforms all baseline retrieval systems, and the improvements lead to better answer generation.
 2 Background
 In this section, we first define the time-sensitive question answering task (§2.1), and then introduce the baseline retrieval-augmented generation QA based systems (§2.2).
-2.1 Time-sensitive Question Answering
-Our task is to predict an answer $y$, to a time-sensitive question $q$, in which the input question $q$ also includes temporal relations $t_{1},...,t_{n}$.
-[^1] To simplify our setting, we make sure each question has a valid answer.
-With advancements in LLMs, a straightforward approach is to directly ask these questions to an off-the-shelf LLM (e.g., LLama 3.1), hoping the relevant facts are captured in its parametric knowledge.
-However, this is often not the case for time-sensitive questions.
-A significant portion of such questions are recent events (e.g., “Who is the UK Prime Minister in November 2024?”), which typically fall beyond the knowledge cutoff for training LLMs (e.g., December 2023 for LLama 3.1).
-One potential solution is to constantly update the parameters of LLMs with new facts, i.e., knowledge editing (Wu et al., 2024b; Wang et al., 2024).
-However, this approach is highly resource-intensive and often impractical for very large LLMs (e.g., GPT-4o).
-Moreover, knowledge editing is not always effective.
-Recent studies reveal that even after parameter updates, LLMs often fail to accurately answer questions about new facts (Liška et al., 2022; Rozner et al., 2024; Wang et al., 2024), let alone related or associated facts (Zhong et al., 2023; Wu et al., 2024b, c).
-Therefore, our study adopts an alternative approach that leverages external text collections as a knowledge source, integrating relevant information into existing LLMs to answer questions (detailed in §2.2).
-While we use Wikipedia222we use December 2021 Wikipedia dump, with 33.1M text chunks as the text collection $D$, our approach is broadly applicable to any corpus, e.g., ClueWeb (Overwijk et al., 2022).
+2.1 Temporal Question Answering
+There has been extensive research on Temporal Question Answering over Knowledge Graphs (TKGQA).
+Prior works have developed knowledge graph-based benchmarks (e.g., CronQuestions (Saxena et al., 2021) and TimeQuestions (Jia et al., 2021)), graph neural network models (Chen et al., 2022; Liu et al., 2023; Sharma et al., 2023), and TKG-augmented LLMs (Qian et al., 2024).
+However, these methods heavily rely on knowledge graphs, which are costly to build, domain-specific, and prone to becoming outdated (Kau et al., 2024).
+Therefore, we focus on the free-text setting, which is more universal and up-to-date.
+Instead of building temporal-specific graph models, we improve general-purpose text retrievers for temporal queries.
+We use Wikipedia as the main knowledge corpus $D$ [^2] December 2021 Wikipedia dump, comprising 33.1 million text chunks.
+Our approach is broadly applicable to other collections, such as the New York Times Annotated Archive (Sandhaus, 2008) and ClueWeb (Overwijk et al., 2022).
 2.2 Retrieval-Augmented Generation
-The goal of RAG is to address the limitations in the parametric knowledge of LLMs (Izacard et al., 2020).
-RAG involves a retrieval system to find relevant evidence from a large corpus and a LLM to generate a response based on the evidence (Wu et al., 2024a).
+The goal of RAG is to address the limitations in the parametric knowledge of LLMs by incorporating external knowledge.
 Passage retrieval and reranking.
 Retrieval methods are typically categorized into sparse retrieval and dense retrieval.
-Sparse retrieval methods, like TF-IDF (Das and Chakraborty, 2018) and BM25 (MacAvaney et al., 2020), rely on lexical matching.
-In contrast, dense retrieval models (Karpukhin et al., 2020; Zhao et al., 2021; Thakur et al., 2021; Izacard et al., 2021) use LLMs (e.g., BERT) encode the question $q$ and the passage $p$ separately using two independent encoders (i.e., bi-encoders).
-These models learn a scoring function (e.g., dot product) between question and passage vectors to capture semantic similarities:
+Sparse retrieval methods like BM25 (MacAvaney et al., 2020) rely on lexical matching.
+In contrast, dense retrieval models (Karpukhin et al., 2020; Zhao et al., 2021; Thakur et al., 2021; Izacard et al., 2021) encode the question $q$ and passage $p$ into low-dimension vectors separately.
+The semantic similarity is computed using a scoring function (e.g., dot product) as:
 $f(q,p)=\text{sim}(\text{Enc}_{Q}(q),\text{Enc}_{P}(p)),p\in\mathcal{D}.$ (1)
-Dense retrieval models are highly scalable, since passages can be encoded offline, and are efficiently retrieved over maximum inner product search (MIPS) with the question (Shrivastava and Li, 2014).
-However, bi-encoder models may lack the ability to capture fine-grained interactions between the query and passage.
+However, these bi-encoder models lack the ability to capture fine-grained interactions between the query and passage.
 A common optional [^3] Note that reranking is not always adopted, as it adds additional computational cost.
-approach is to have another cross-encoder model to rerank top retrieved passages.
+approach is to have another cross-encoder model to rerank top passages.
 Cross-encoder models (Khattab and Zaharia, 2020; Wang et al., 2020; Gemma et al., 2024) jointly encode the query $q$ and the passage $p$ together by concatenating them as input into a single model as:
@@
 Answer generation.
-The answer generation component takes the question and support passages as input to generate the answer.
-A widely used approach is Fusion-in-Decoder (Izacard et al., 2020), which fine-tunes an encoder-decoder LLM (e.g., T5 (Raffel et al., 2020)).
-FiD independently encodes each passage and question, then concatenates their representations into the decoder, as formulated below:
-$y=\text{Dec}([\text{Enc}([q;p_{1}]);\dots;\text{Enc}([q;p_{k}])]),p_{k}\in\mathcal{D}.$ (3)
-With decoder-only LLMs (OpenAI et al., 2024; Dubey et al., 2024), the support passages and the question are concatenated into a single sequence.
-To unlock the reasoning capabilities of LLMs, Chain-of-Thought (CoT) prompting (Wei et al., 2022) introduces intermediate reasoning steps within prompt examples, significantly improving performance on various reasoning tasks.
-Self-RAG (Asai et al., 2024) predicts answers for each passage independently, and then adopts self-reflect mechanisms (Yao et al., 2022) into prompts to select the best answer.
-Self-RAG achieve SOTA results on multiple knowledge-intensive QA datasets.
+Recent reader systems are mainly powered by LLMs with strong reasoning capabilities.
+With recent advancements in long-context LLMs (Dubey et al., 2024; Lee et al., 2024b), top documents are typically concatenated with the query as reader input:
+$y=\text{Dec}(\text{Enc}([p_{1};\dots;p_{k};q])).$ (3)
+To unlock the reasoning capabilities of LLMs, Chain-of-Thought (CoT) prompting (Wei et al., 2022) introduces intermediate reasoning steps for improved performance.
+Self-RAG (Asai et al., 2024) critiques the retrieved passages and its own generations.
+Recent agentic RAG systems dynamically acquire and integrate external knowledge (using either retriever models or search APIs) during the reasoning process (Li et al., 2025; Song et al., 2025; Jin et al., 2025).
 3 TempRAGEval Benchmark
 In this section, we first present existing time-sensitive QA datasets (§3.1), then introduce our diagnostic benchmark dataset TempRAGEval3.2), finally we evaluate existing retrieval approaches on TempRAGEval3.3).
-3.1 Existing Time-Sensitive QA Datasets
-Existing time-sensitive QA datasets often feature questions where timestamps are closely tied to their answers, as illustrated in Figure 1.
-Consequently, retrievers relying on surface-form date matching tend to perform well on these test sets, limiting the evaluation of true temporal reasoning capabilities.
-To systematically evaluate time-sensitive question answering, we establish the following criteria: (1) questions with varying difficulty levels, including both common and rare entities; (2) manual evidence annotation to enable retrieval evaluation; and (3) temporal perturbations as contrast sets (Gardner et al., 2020).
-Since no existing datasets provide both temporal perturbations and evidence annotations, we primarily adapt from the following two datasets:
-  • SituatedQA (Zhang and Choi, 2021) is a open-domain QA dataset where the correct answer to a question is based on a temporal context, specifically of the type “as of”.
-SituatedQA sources questions from exiting QA datasets such as NQ-Open (Kwiatkowski et al., 2019) and WebQuestions (Berant et al., 2013), which primarily focus on popular entities.
-  • TimeQA (Chen et al., 2021) is a time-sensitive QA dataset in which question-answer pairs are synthetically generated from time-evolving WikiData facts using templates.
-The dataset includes a wide range of reasoning types and difficulty levels, with a notable emphasis on rare entities.
-Other time-sensitive QA datasets, StreamingQA Liška et al.
-(2022), TempLAMA (Dhingra et al., 2022), and concurrent dataset ComplexTQA (Gruber et al., 2024), can also serve as alternatives.
-Temporal reasoning benchmarks such as Wei et al.
-(2023); Chu et al.
-(2024); Wang and Zhao (2024); Tan et al.
-(2023) fall outside the scope of our study.
-3.2 TempRAGEval Construction
-Dataset # Eval. Evid. Natu. Pert.
-ComplexTQA 10M
+Dataset # Eval. Evid. Natu. Comp.
+ComplexTQA 10M $\checkmark$
 StreamingQA 40K $\checkmark$ $\checkmark$
@@
 SituatedQA 2K $\checkmark$
-TimeQA 3K
-MenatQA 2K
+TimeQA 3K $\checkmark$
+MenatQA 2K $\checkmark$
 TempRAGEval 1K $\checkmark$ $\checkmark$ $\checkmark$
-Table 1: Comparison of temporal QA datasets. TempRAGEval is featured by manual evidence annotations, human-written question (i.e., Naturalness), and perturbations.
+Table 1: Comparison of temporal QA datasets. TempRAGEval is featured by manual evidence annotations, human-written question (i.e., Naturalness), and higher complexity in temporal reasoning.
+3.1 Existing Time-Sensitive QA Datasets
+There are several existing QA datasets that focus on temporal reasoning.
+The most representative ones are the following: [^4] We mainly focus on these two datasets, while others, such as Liška et al.
+(2022); Dhingra et al.
+(2022); Gruber et al.
+(2024), serve as alternatives.
+Datasets such as Tan et al.
+(2023); Virgo et al.
+(2022) that are not knowledge-intensive, are excluded from this work (Appendix B).
+  • SituatedQA (Zhang and Choi, 2021) is a time-sensitive QA dataset where the answer to an information-seeking question varies based on temporal context.
+These questions contain a single type of temporal constraint (e.g.,as of”) that directly align with the answers.
+Retrievers with surface-form date matching often exploit these shortcuts to bypass the need for temporal reasoning.
+  • TimeQA (Chen et al., 2021) is another time-sensitive QA dataset.
+Unlike SituatedQA, the questions in hard split include complex temporal constraints (e.g.,between 2012 to 2018”).
+However, question-answer pairs are synthetically generated from time-evolving WikiData facts using templates.
+In addition, TimeQA does not include evidence annotations, making it imprecise to evaluate retrieval results.
+According to Table 1, we observe that none of existing datasets include these key factors for systematically evaluating current retrieval (and answer generation) systems: (1) Evidence annotation; (2) Natural questions from users; (3) Complex temporal constraints.
+Therefore, as we will show in the following section, we aim to address this limitation by creating TempRAGEval.
+3.2 TempRAGEval Construction
 We create TempRAGEval, a time-sensitive QA benchmark for rigorously evaluating temporal reasoning in both retrieval and answer generation.
 Perturbed question-answer pair generation.
-We source questions from TimeQA (Chen et al., 2021) and SituatedQA (Zhang and Choi, 2021) datasets.
-For each question, we introduce perturbations to the original temporal constraints, e.g.,as of 2019”, with complex variations such as e.g.,between 2020 and 2021”, “the last before 2020”, or “as of 6 May 2021”.
-We guarantee the answer (e.g.,Boris Johnson”) remains consistent across the perturbed questions.
+Annotators first select question-answer pairs from the SituatedQA and TimeQA datasets [^5] Since both datasets lack questions about knowledge beyond the cutoff date of existing LLMs, we primarily focus on historical knowledge and discuss potential future directions on recent knowledge in the Limitations section.
+that can be grounded in Wikipedia facts with key timestamps or durations.
+They then revise temporal perturbations by selecting implicit conditions, temporal relations, and alternative dates to include complex temporal reasoning, without changing the final answer [^6] Questions with the same content but different temporal constraints and answers are considered different samples.
+Perturbations are introduced for each sample..
+Annotators are also required to edit the question text to improve naturalness.
+We include detailed guidelines in Appendix C.
 Evidence annotation.
 To better evaluate the performance of retrieval systems, we supplement question-answer pairs with up to two annotated gold evidence passages.
-Specifically, for each question, we manually review top-100 passages retrieved by the leading retriever, Contriever (Izacard et al., 2021), and the best reranker, Gemma (Li et al., 2023).
-If these retrieval methods fail to identify relevant passages within top-100 passages, we manually search Wikipedia pages related to the query entities to locate the gold evidence (around 12.7% of original questions).
-In total, we have annotated 1,000 test examples with both original and perturbed questions, along with human-annotated evidence.
-TempRAGEval-TimeQA (500 examples) focus on popular entities, and TempRAGEval-TimeQA (500 examples) on rare entities.
+A passage is relevant to the question if annotators can obtain the correct answer based on the passage.
+Specifically, for each question, annotators are asked to manually review top-20 passages retrieved by Contriever (Izacard et al., 2021) and reranked by the best Gemma (Li et al., 2023) reranker.
+If there is no relevant passage, annotators are required to search Wikipedia pages related to the query entities to locate the gold evidence (around 12.7% of questions).
+We create 1,000 test examples with human-annotated evidence.
+Appendix D presents sample statistics, revealing that SituatedQA questions include popular entities while the entities for TimeQA questions are long-tailed.
 3.3 Preliminary Evaluation on TempRAGEval
@@
 We present full results in Figure 5 in Appendix.
-Figure 2: The retrieval performance degradation of the Gemma baseline on TempRAGEval-SituatedQA, comparing original and perturbed questions (see TempRAGEval-TimeQA in Appendix C).
+Figure 2: The retrieval performance degradation of the Gemma baseline on TempRAGEval-SituatedQA, comparing original and perturbed questions (see TempRAGEval-TimeQA in Appendix E).
 4 MRAG: Modular Retrieval
@@
 Retrieval and summarization.
-We apply off-the-shelf retrievers (e.g., Contriever) to find relevant passages to MC in Wikipedia. Then we adopt a cross-encoder model to rerank the passages with high semantic similarity to MC.
+We apply off-the-shelf retrievers (e.g., Contriever) to find relevant passages to MC in Wikipedia. Then we employ reranker models to reorder these passages by semantic similarity to MC.
 It is common for a passage to contain multiple pieces of temporal information, most of which are unrelated to the question and can distract the temporal scoring component introduced next.
 For example, the relevant passage in Figure 3 includes the sentence, “it was filmed in 2017”, which satisfies the TC but is irrelevant.
-Therefore, we initially considered splitting passages into individual sentences to eliminate temporal distractors.
-However, as shown in Figure 3, critical information from the most relevant passages—such as “America’s Next Top Model”, “The winner of the competition”, and ”January 9, 2018”—is often scattered across different sentences.
+Therefore, we split passages into individual sentences to eliminate temporal distractors.
+However, as shown in Figure 3, critical information from the most relevant passages—such as “America’s Next Top Model”, “The winner of the competition”, and ”January 9, 2018”—can be scattered across different sentences.
 Relying solely on sentence splitting would miss key details.
-To overcome this challenge, we instead adopt LLMs to summarize the top-k passages into a single sentence to condense relevant phrases and temporal information.
-[^4] While using LLMs to summarize passages can introduce hallucinations, we mitigate this by summarizing only the top-k passages and simply splitting the remaining passages into sentences.
-An analysis of this approach is provided in §6.1.
+To overcome this challenge, we additionally employ LLMs to summarize each of the top-k passages into a single sentence condensing relevant phrases and temporal information, as analyzed in §6.1.777While using LLMs to summarize passages can introduce hallucinations, we mitigate this by summarizing only the top-k passages.
 Semantic-temporal hybrid ranking.
@@
 The semantic score is calculated from the similarity between the evidence sentence and MC.
-For temporal score, we first extract temporal information in the evidence sentence, then we apply symbolic scoring functions [^5] We include details of pre-defined temporal scoring functions in Appendix D.
-to verify whether the temporal information in the evidence satisfies TC.
-For example, the evidence summary timestamp “2018” satisfies TC “as of 2021” in Figure 3.
-The final score for each sentence is the product of the semantic score and the temporal score.
-We select the passages that contains the highest-scoring sentences.
+For temporal score, we first extract the timestamp from each sentence (e.g.,2018”).
+Based on the timestamp and TC (e.g.,as of 2021”), we compute a temporal score using symbolic functions similar to temporal activation functions in Chen et al.
+(2022).
+The final score for each sentence is obtained by multiplying the semantic score and the temporal score.
+Finally, we select the passages that contains the highest-scoring sentences.
+We include the details of symbolic functions in Appendix G.
 5 Experiments
@@
 For retrieval, we include BM25, Contriever, and a hybrid method (Jedidi et al., 2024).
-Reranking methods include ELECTRA (Clark et al., 2020), MiniLM (Wang et al., 2020), Jina(Jina, 2024), BGE (Xiao et al., 2023), NV-Embed (Lee et al., 2024), and Gemma (Gemma et al., 2024).
+Reranking methods include ELECTRA (Clark et al., 2020), MiniLM (Wang et al., 2020), Jina(Jina, 2024), BGE (Xiao et al., 2023), NV-Embed (Lee et al., 2024a), and Gemma (Gemma et al., 2024).
 We follow state-of-the-art answer generation approaches based on prompting LLMs (§2.2).
@@
 We use 10 examples in prompts.
+To eliminate the impact of input length constraints across models, we conduct parametric studies as described in LABEL:sec:_concat and report each LLM’s performance with its optimal number of input passages in Table 3.
 Method TempRAGEval-TimeQA TempRAGEval-SituatedQA
@@
 Bold numbers indicate the best performance.
-We include complete results in Appendix E.
+We include complete results in Appendix I.
 5.2 Main Results
@@
 Self-MRAG 5 61.2 75.3 20 72.2 76.0
-Table 3: End-to-end QA comparison for various generation strategies and LLMs on TempRAGEval. Bold numbers indicate the best performance the each backbone LLM. The second best is underlined.
+Table 3: End-to-end QA performance comparison for various generation strategies and LLMs on TempRAGEval.
+Bold numbers indicate the best performance the each backbone LLM.
+The second best is underlined.
+TIMO has a limited input length with up to three passages.
+We report the best number of passages for Llama models, and provide ablation on different numbers in LABEL:sec:_concat.
 MRAG enhances retrieval performance for time-sensitive questions.
@@
 As shown in Table 3, MRAG outperforms baseline RAG approaches in QA accuracy, for instance 49.2% EM (MRAG) over 44.0% EM (RAG) in TempRAGEval-TimeQA for Llama3.1-8B.
-This highlights the role of enhanced retrieval in improving accuracy of the time-sensitive QA task.
-Incorporating a self-reflection strategy improves performance for Llama3.1 models but not for TIMO.
-We observe that TIMO makes more errors in self-reflection, likely due to the limited reasoning capacity of its backbone model, Llama 2.
+Incorporating a self-reflection strategy improves performance for Llama3.1 models but not for TIMO, likely due to the limited reasoning capacity of its backbone model, Llama 2.
 6 Analysis
-In this part, we provide detailed analysis on how each component of MRAG contributes to the overall improvement.
+This section presents a detailed analysis of the results and the contribution of each MRAG module.
 6.1 Ablation Study
-The impact of the numbers of passages for summarization.
-Our LLM based summarization removes irrelevant temporal information but may also introduce hallucinations.
-We conduct a parametric study to evaluate the impact of the number of passages used for summarization.
-As shown at the bottom of Table 2, summarizing top-five passages achieves the best balance.
-The impact of the number of passages used for answer generation in MRAG.
-Our experiments show that the optimal number of passages depends on the size of the LLM.
+The impact of the number of passages for summarization.
+Our LLM based summarization removes irrelevant temporal information but may also introduce hallucinations (details in Appendix H).
+The parametric study at the bottom of Table 2 shows that summarizing top-five passages achieves the best balance.
+Additionally, we compare the RAG setup with the long-document QA setup bypassing retrievers in LABEL:longdoc.
+RAG achieves a better accuracy, as retrievers exclude irrelevant passages and reduce noise.
+The impact of the number of passages for answer generation.
+Our experiments show that the optimal number of passages depends on the LLMs.
 TIMO can handle a maximum of 3 passages, while the optimal number for LLaMA 3.1-8B is five, and for LLaMA 3.1-70B, it is twenty.
-Larger LLMs, with more parameters, process longer contexts better.
 Full results are presented in LABEL:sec:_concat.
+The impact of key retrieval steps in MRAG.
+In the retrieval and summarization module, with retrieved passages, MRAG conducts two key steps: (1) Passage Keyword Ranking reduces the passages from 1,000 to 100 based on the keyword presence; (2) Passage Semantic Ranking reorders the top-100 passages and splits them into $\sim$500 sentences, including chunk summaries.
+Similarly, in the semantic-temporal hybrid ranking module, another two key steps are conducted: (1) Sentence Keyword Ranking further narrows the scope to 200 sentences; (2) Hybrid Ranking ranks these sentences by semantic-temporal hybrid scoring.
+As shown in LABEL:ablation, keyword-based ranking steps effectively reduce ranking candidates without significant performance loss, while semantic and hybrid ranking steps markedly enhance retrieval performance.
+An efficiency-accuracy trade-off can be made by adjusting the number of targeted passages or sentences at each step.
+Computational Overhead.
+As MRAG involves retrieval, summarization, and re-ranking, it incurs approximately twice the computational overhead of standard RAG pipelines, which is manageable.
+We provide a detailed analysis in LABEL:overhead.
 6.2 Human Evaluation
@@
 To address this, we conduct a human evaluation on a random subset of 200 examples to assess actual retrieval performance.
-We manually review the top 20 passages retrieved by MRAG and Gemma, annotating the first relevant passage.
-The results confirm that MRAG consistently outperforms Gemma in retrieval accuracy.
-We present full results in LABEL:sec:_human.
+The results validate the advantages of MRAG over Gemma in retrieval accuracy with full results presented in LABEL:sec:_human.
 6.3 Case Study
-Figure 4: A case study for top-1 passages retrieved by Gemma and MRAG for a question from TempRAGEval-SituatedQA.
-We conduct case studies to qualitatively assess the advantages of MRAG over the Gemma baseline.
-According to Figure 4, the top-1 passage by Gemma matches the query date “1988”, but discusses a father-son record set in 2007.
-In contrast, MRAG result is about a teammate combination record in the same season, despite the date “1961” differing from the query date “1988”.
-We include more cases in LABEL:sec:_cases.
+Figure 4: A case study for top-1 passage retrieved by Gemma and MRAG from TempRAGEval.
+We analyze retrieval errors qualitatively to highlight the advantages of MRAG over the Gemma baseline.
+As shown in Figure 4, the top-1 passage by Gemma matches the query date “1988” but discusses a father-son record set in 2007.
+In contrast, MRAG retrieves a passage about a teammate combination record from the same season, despite the differing date (“1961” vs.
+“1988”).
+Additional cases and answer generation error cases are provided in LABEL:sec:_cases_retrieval and LABEL:sec:_cases_qa.
 7 Other Related Works
@@
 Some studies focus on distilling or fine-tuning LLM embeddings for reranking tasks, such as Gemma (Gemma et al., 2024) and MiniCPM (Hu et al., 2024).
-Others aim to develop generalist embedding models capable of performing a wide range of tasks including retrieval and reranking, e.g., gte-Qwen (Yang et al., 2024) and NV-Embed (Lee et al., 2024).
-These LLM-based methods have demonstrated unprecedented performance in benchmarks such as MTEB (Muennighoff et al., 2023) over Contriver-based dense retrievers, and also on our TempRAGEval benchmark.
+Others aim to develop generalist embedding models capable of performing a wide range of tasks including retrieval and reranking, e.g., gte-Qwen (Yang et al., 2024) and NV-Embed (Lee et al., 2024a).
+These LLM-based methods have demonstrated unprecedented performance in benchmarks such as MTEB (Muennighoff et al., 2023) and our TempRAGEval.
 Reasoning intensive retrieval.
@@
 Su et al.
-(2024a) introduces BRIGHT, a new retrieval benchmark emphasizing intensive reasoning.
-The benchmark challenges retrievers to perform reasoning across diverse domains, such as analogies and algorithms, to identify relevant documents.
-We focus on temporal reasoning-intensive retrieval, sharing a similar vision for future advanced retrieval systems.
+(2024a) introduces BRIGHT, a new retrieval task emphasizing intensive reasoning.
+We focus on temporal reasoning, one aspect of a broader class of reasoning-intensive retrieval.
+MRAG is expected to generalize to other forms of symbolic reasoning, such as numeric ranges and geospatial constraints.
+It mitigates direct numeric matching in retrieval and enhances reasoning capabilities.
 8 Conclusion
 This study focuses on time-sensitive QA, a task that challenges LLM based QA systems.
-We first present TempRAGEval, a diagnostic benchmark featuring natural questions, evidence annotations, and temporal perturbations.
+We first present TempRAGEval, a diagnostic benchmark featuring natural questions, evidence annotations, and temporal complexity.
 We further propose a training-free MRAG framework, which disentangles relevance-based retrieval from temporal reasoning and introduces a symbolic temporal scoring mechanism.
@@
 However, addressing questions with implicit temporal constraints presents a more complex challenge for retrieval systems.
-Our symbolic approach is not directly applicable to such questions and requires more advanced reasoning mechanisms to bridge this gap.
-(2) Our benchmark primarily targets time-sensitive questions with a single answer.
-Future work can extend to predict a list of answers with different temporal constraints.
-(3) Despite our approach improves temporal reasoning in retrieval, there is still head room.
-Fine-tuning dense retrieval models on temporal reasoning-intensive retrieval datasets may offer a promising solution for future work.
+We could extend to implicit temporal reasoning by associating explicit information to the implicit one using LLM common sense and background knowledge like (Chen et al., 2022).
+(2) Our dataset does not include time-sensitive questions that fall outside the LLM knowledge cutoff.
+We could extend our dataset with questions from RealTime QA (Kasai et al., 2024) and AntiLeak-Bench (Wu et al., 2024b).
+(3) Our main objective is to improve temporal reasoning in retrieval, which has not been tackled by previous works.
+More complex scenarios like multi-hop and recursive reasoning require further research efforts.
+(4) The proposed framework introduces computational overhead for improved performance as detailed in LABEL:overhead.
+(5) We analyze knowledge conflicts between LLMs and passages in LABEL:sec:_conflict and leave conflicts among passages for future work.
+Ethics Statement
+TempRAGEval were constructed upon the test set of TimeQA (Chen et al., 2021) and SituatedQA (Zhang and Choi, 2021) datasets, which are publicly available under the licenses BSD-3-Clause license [^8] https://opensource.org/licenses/BSD-3-Clause and Apache-2.0 license [^9] https://www.apache.org/licenses/LICENSE-2.0.
+These licenses all permit us to compose, modify, publish, and distribute additional annotations upon the original dataset.
+All the experiments in this paper were conducted on 4 NVIDIA L40S 46G GPUs.
+We hired 3 graduate students in STEM majors as annotators.
+We recommended that annotators spend at most 2 hours per day for annotation in order to reduce pressure and maintain a comfortable pace.
+The whole annotation work lasted about 5 days.
 Appendix A Controlled Experiments
@@
 However, for “before” and “as of”, despite their temporal relevance, documents with earlier dates fail to achieve sufficiently high similarity scores, potentially leading to their exclusion from the retrieval process.
-In summary, existing retrieval methods demonstrate limited temporal reasoning capabilities. The LLM embedding-based method shows better performance than others.
+In summary, existing retrieval methods demonstrate limited temporal reasoning capabilities.
+The LLM embedding-based method shows better performance than others.
+Our proposed MRAG framework is retriever-agnostic, which aims to improve temporal reasoning capabilities for any type of retrieval models.
 Figure 5: Similarity scores of query-document pairs by varying the temporal relation in the query and the date in the document.
-Appendix B Implementation Details
+Appendix B Dataset Selection Criteria
+Our benchmark focuses on time-sensitive question answering, which is knowledge-intensive.
+Therefore, datasets designed for only temporal reasoning (e.g., “What is the time 5 year and 5 month after Oct, 1444”) are not considered, such as TempReason (Tan et al., 2023) and DurationQA (Virgo et al., 2022).
+Aggregated benchmarks (e.g., TimeBench (Chu et al., 2024) and TRAM (Wang and Zhao, 2024)) focus on evaluating diverse temporal reasoning capabilities, which have a broader scope than our focus.
+MenatQA (Wei et al., 2023) is built by adding counterfactual and order factors to TimeQA (Chen et al., 2021) questions, which is similar to our approach.
+Other knowledge-intensive temporal QA datasets can serve as alternative sample sources including StreamingQA Liška et al.
+(2022), TempLAMA (Dhingra et al., 2022), and concurrent dataset ComplexTQA (Gruber et al., 2024).
+We select SituatedQA (Zhang and Choi, 2021) for its human-written questions, distinguishing it from other temporal QA datasets that are typically synthetic.
+Additionally, we opt for TimeQA (Chen et al., 2021) due to its hard split, which already includes complex temporal questions.
+Notably, both SituatedQA and TimeQA can be grounded in the Wikipedia corpus.
+Appendix C Annotation Guidelines
+C.1 Annotating Perturbations
+Given a question-answer pair sourced from TimeQA or SituatedQA (e.g., Q: “Arnolfini Portrait was owned by whom between Jul 1842 and Nov 1842?” A: “National Gallery”), annotators should ground the pair to facts in Wikipedia (e.g., “The Arnolfini Wedding by Jan van Eyck, has been part of the National Gallery’s collection in London since 1842.”).
+Then they identify the key timestamps or durations of Wikipedia facts (e.g., 1842).
+To create temporal perturbations, annotators are asked to come up with combinations of implicit conditions, temporal relations, and alternative dates to form complex temporal constraints.
+The implicit condition can be “None” or selected from a list of 4 types: “first”, “earliest”, “last”, and “latest”.
+The temporal relation should be selected from a list of 11 types: “as of”, “from to”, “until”, “before”, “after”, “around”, “between”, “by”, “in”, “on”, and “since”.
+Finally, annotators rewrite questions naturally (e.g., “Who is the last one owned Arnolfini Portrait after 1700?”) by introducing perturbed temporal constraints (e.g., “last … after 1700”) and ensure that the answers (e.g., “National Gallery”) remain unchanged.
+After different annotators create perturbed question-answer pairs, they exchange these pairs with each other to validate the correctness of the answers.
+Only the perturbed samples validated by two annotators are kept.
+C.2 Annotating Gold Evidences
+For gold evidence annotations, annotators are assigned different perturbed question-answer pairs.
+For each pair, 20 context passages are provided to annotators, which are retrieved by the leading retriever Contriever (Izacard et al., 2021), and the best reranker Gemma (Li et al., 2023).
+Annotators are asked to identify up to two gold evidence passages from these passages.
+A passage is regarded as relevant and annotated as gold evidence if annotators can obtain the correct answer from this passage.
+If there is no relevant passage among these 20 retrieved ones, annotators should search Wikipedia pages related to the query entities to locate the gold evidence passages manually.
+Lastly, annotators exchange samples to validate gold evidence annotated by others.
+Only the gold evidence annotations validated by two annotators are kept.
+Appendix D Sample Statistics
+We gather a similar size of examples as previous temporal QA benchmarks (e.g., 3K for TimeQA (Chen et al., 2021) and 2K for MenatQA (Wei et al., 2023)), which is enough for an evaluation set (see examples in LABEL:examples).
+As we manually annotate gold evidence passages in Wikipedia, it is time-consuming to scale up like other synthetic datasets (Dhingra et al., 2022; Gruber et al., 2024).
+To understand the difference between two subsets, we summarize the statistics in Appendix D.
+The average length of questions is measured by the GPT-2 tokenizer (Radford et al., 2019).
+We assess the popularity of key entities in questions using the average monthly page view counts of the corresponding Wikipedia page in 2024 (Pageviews, 2024).
+As we can see, the main difference lies in the question entity popularity.
+TempRAGEval-TimeQA questions typically inquire about lesser-known individuals and are generally straightforward and clear.
+In contrast, TempRAGEval-SituatedQA questions commonly ask about a sports team and championship, which are more general and sometimes ambiguous.
+This difference may explain varying retrieval and QA performance across the two subsets.
+TimeQA SituatedQA
+# original questions 123 120
+# perturbed questions 377 380
+# total questions 500 500
+Temporal complexity hard hard
+Avg. question length 15.2 15.6
+Avg. entity popularity 7,456 57,521
+Table 4: Sample statistics for TempRAGEval-TimeQA and TempRAGEval-SituatedQA.
+Figure 6: Retrieval performance difference between original queries and perturbed queries in TempRAGEval subsets for the baseline Gemma retrieval.
+Appendix E Retrieval Performance Degradation Due to Perturbations
+As shown in Figure 6, we compare the retrieval performance between original queries and perturbed queries using the same baseline retrieval system (i.e., Contriever retriever and Gemma reranker).
+For both the TimeQA and SituatedQA subsets, the perturbed questions significantly increase the difficulty of retrieving relevant documents, particularly when evaluating the top-1 and top-5 ranked documents.
+This suggests that the introduction of perturbations introduces greater complexity.
+The existing retrieval method has limited temporal reasoning capabilities and is not robust to such variations.
+Appendix F Evaluation Experiment Implementation Details
 We conduct empirical evaluations for MRAG and SOTA retrieving-and-reranking systems on TempRAGEval.
+In baselines, due to limited computing resources, we use LLM-based embedding models as a reranking model, such as Gemma (Gemma et al., 2024) and NV-Embed (Lee et al., 2024a).
 MRAG consists of functional modules, which can be based on algorithms, models, or prompting methods.
@@
 LLM prompting based modules are keyword extraction, query-focused summarization.
-As shown in Appendix E, we have tested Llama3.1-8B-Instruct and Llama3.1-70B-Instruct models (Dubey et al., 2024) for LLM prompting based modules.
+As shown in Appendix I, we have tested Llama3.1-8B-Instruct and Llama3.1-70B-Instruct models (Dubey et al., 2024) for LLM prompting based modules.
 Detailed prompts are listed in LABEL:sec:_prompts.
 The evaluation metrics are computed based on retrieved passages not sentences.
-Appendix C Retrieval Performance Degradation from Perturbations
-We compare the retrieval performance between original queries and perturbed queries using the same baseline retrieval system (i.e., Contriever retriever and Gemma reranker).
-For both the TimeQA and SituatedQA subsets, the perturbed questions significantly increase the difficulty of retrieving relevant documents, particularly when evaluating the top-1 and top-5 ranked documents.
-This suggests that the introduction of perturbations introduces greater complexity.
-The existing retrieval method has limited temporal reasoning capabilities and is not robust to such variations.
-Figure 6: Retrieval performance difference between original queries and perturbed queries in TempRAGEval subsets for the baseline Gemma retrieval.
-Appendix D Spline Functions for Temporal Scoring
-In MRAG, the temporal relevance is determined symbolically using temporal constraints in query, extracted dates from sentences, and pre-defined scoring functions.
-The temporal constraints are categorized by implicit conditions, i.e., “first” and “last”, and temporal relations, i.e., “before”, “after”, and “between”.
-For example, the question “Who won the latest game as of 1981?” will be classified as the “last” type and the “before” type.
-We pre-define spline function templates for each type of temporal constraints, as shown in Figure 7.
-Each spline function computes a temporal score for a given input date extracted from the sentence.
-For sentences with multiple dates, the highest temporal score is selected.
 Figure 7: Pre-defined spline functions for temporal relevance scoring.
@@
 The horizontal coordinate of each subplot is the date in the document sentence.
-Appendix E Complete Retrieval Evaluation Results
+Appendix G Implementations for Semantic-Temporal Hybrid Scoring
+The Retrieval and Summarization module in MRAG splits and summarizes top relevant passages into independent sentences for the downstream fine-grained reranking, which is inspired by Yang (2023).
+The Semantic-Temporal Hybrid Scoring module is designed to assess the semantic relevance and temporal relevance between the question and the evidence sentence.
+To quantify the semantic relevance, we apply an embedding model (e.g., Gemma) to the question main content and the sentence.
+For temporal relevance, we employ a symbolic scoring approach, wherein the module automatically generates scoring functions for each question and computes temporal scores for individual sentences.
+To generate scoring functions, we classify question temporal constraints into six categories and define a template for each.
+A scoring function is instantiated using the corresponding template and extracted timestamp(s) from the question.
+The six constraint types include: “first - before”, “first - after”, “first - between”, “last - before”, “last - after”, and “last - between”.
+Here, “last” denotes that the question seeks the most recent event, while “before” indicates that the event must precede a specified date.
+For instance, the question “Who won the latest game as of 1981?” corresponds to “last - before - 1981”, as illustrated in the top-left subplot of Figure 7.
+Afterwards, the module extracts the timestamp(s) from the evidence sentence.
+It computes the temporal score based on the extracted timestamp and the corresponding scoring function.
+If multiple timestamps are present, the highest temporal score is selected.
+For example, as depicted in Figure 7, for the constraint “last - before - 1981”, an evidence sentence mentioning “1970” would receive a temporal score around 0.9.
+The final score of each evidence sentence for each question is obtained by multiplying the temporal score and the semantic score.
+Finally, we select the passages that contains the highest-scoring sentences.
+The passages are fed into the later answer generation stage rather than the sentences.
+The passages provide better background information, which leads to higher generation quality for reader systems.
+Appendix H LLM-based Summarization Case Study
+LLM-based query-focused summarization enhances retrieval performance by distilling key information from passages while filtering out irrelevant context, as demonstrated in Table 5.
+In the second success case, the summarization effectively converts structured data into natural language, benefiting retrievers that are primarily trained on free-text retrieval tasks.
+However, LLM-generated summaries may introduce hallucinations and errors, though their occurrence is infrequent.
+As shown in Table 5, erroneous summaries can mislead the retriever with non-factual events or incorrect dates, resulting in irrelevant passages ranking higher.
+To balance retrieval improvements with the risk of errors, we summarize only the top-$k$ passages per query, which also reduces computational overhead.
+Furthermore, to prevent error propagation, we provide the reader model with original passages rather than their summaries.
+Success Cases
+Question Who won the latest America’s Next Top Model by May 8, 2021?
+Answer Kyla Coleman
+Passage America’s Next Top Model (season 24) | The twenty-fourth cycle of America’s Next Top Model premiered on January 9, 2018 … The winner of the competition was 20 year-old Kyla Coleman from Lacey, Washington with Jeana Turner placing as the runner up.
+Summarization Kyla Coleman, a 20-year-old from Lacey, Washington, won the competition in 2018.
+Question When was the last time Kentucky won NCAA in basketball after 2010?
+Answer 2012
+Passage Kentucky Wildcats | Men (8) ; Basketball (8): 1948, 1949, 1951, 1958, 1978, 1996, 1998, 2012 ; Women (2) … List of NCAA schools with the most NCAA Division … Kentucky has won 13 NCAA team national championships.
+Summarization The Kentucky Wildcats won the NCAA basketball championship in 1948, 1949, 1951, 1958, 1978, 1996, 1998, and 2012.
+Failure Cases
+Question When was the last time the Ducks won the Stanley Cup as of 2010?
+Answer 2007
+Passage Anaheim Ducks | … Despite the arenas being six hours away from each other, the teams have developed a strong rivalry, primarily from the 2009 and 2018 Stanley Cup playoffs.
+The Ducks won the series in 2009, but the Sharks came back in 2018.
+Summarization The Anaheim Ducks won the Stanley Cup in 2009.
+True fact The Anaheim Ducks won the Stanley Cup in 2007. That was their first and only championship so far.
+Question How many times has South Korea held the Winter Olympics as of 2018?
+Answer 1 | one
+Passage 2018 Winter Olympics | The 2018 Winter Olympics … This marked the second time that South Korea had hosted the Olympic Games (having previously hosted the 1988 Summer Olympics in Seoul) …
+Summarization South Korea held the Winter Olympics in 2018 and previously in 1988.
+True fact South Korea held the Winter Olympics in 2018 and Summer Olympics in 1988.
+Table 5: Success and error cases of LLM-based query-focused summarization using Llama3.1-8B-Instruct.
+Appendix I Complete Retrieval Evaluation Results
+We evaluate MRAG on TempRAGEval with baseline retrieval methods, including ELECTRA101010cross-encoder/ms-marco-electra-base, MiniLM 111111cross-encoder/ms-marco-MiniLM-L-12-v2, Jina121212jinaai/jina-reranker-v2-base-multilingual, BGE [^13] BAAI/bge-reranker-large, NV-Embed141414nvidia/NV-Embed-v1, and Gemma  [^15] BAAI/bge-reranker-v2-gemma.
+Complete rsults are presented in LABEL:experiments_timeqa and LABEL:experiments_situatedqa.